OpenStudy (anonymous):

Solve the following word problem using two variables. Show each step of your work. Tanya is five years younger than Heidi. The sum of their ages is 23. Find both ages.

OpenStudy (anonymous):

@chucho78 @KamiBug

OpenStudy (here_to_help15):

Let the age of Tanya be x and age of heidi be y y = x+5 x+ y = 23 do the substitution method

OpenStudy (here_to_help15):

x + x + 5 = 23

OpenStudy (kamibug):

You can use this system of equations: t = h - 5 t + h = 23

OpenStudy (here_to_help15):

Or... what @KamiBug says :D

OpenStudy (anonymous):

Still don't really understand...

OpenStudy (kamibug):

To solve for t and h, use the substitution method. Since t = h - 5, you can substitute h - 5 for t in the second equation to solve for h. Then, once you know what h is, substitute the value of h into the first equation to solve for t. :) Here are all the steps... h - 5 + h = 23 2h = 28 h = 14 <--- t = 14 - 5 t = ? <---

OpenStudy (kamibug):

Btw, love the whole username and profile pic thing Lol ^.^ xD

OpenStudy (here_to_help15):

lol ;D

OpenStudy (anonymous):

I got it, thank you :) I used a different equation but it won't let me post it (The post button disappears and I can't scroll down) But I got it, Tanya was 9 and Heidi was 14. Thank you both

OpenStudy (kamibug):

I hate it when that happens with the button. So annoying. -.- Great job, though! :D